The vertices of a polygon are (9, 27), (15, 27), (18, 12), and (6, 12). Find the area of the polygon.
Answer by ikleyn(52778)

.
Even with "not-armed" eye, it is clearly seen from the coordinates, that the polygon is a trapezoid with horizontal bases

of the lengths of 15-9 = 6  and  18-6 = 12 units.


The altitude is 27-12 = 15 units.


The area of a trapezoid is the product of the half of the sum of its bases (which is (6+12)/2 = 9 units) and the altitude,


so the area is 15*9 = 135 square units.

Answer.   The area of the polygon is 135 square units.